OpenStudy (anonymous):

What is the equation of the tangent to each curve at this indicated value. f(x) = (x²-3)(x²+1); x=4

OpenStudy (anonymous):

Do you know how to find the derivative of the function?

OpenStudy (anonymous):

yes, is that all i do ?

OpenStudy (anonymous):

Yes, find the derivative and then substitute the x-value you are given to find the slope at x=4. Then just use that as a slope and use it to find the equation of the line.

OpenStudy (anonymous):

Ohhh ! thanks :D

OpenStudy (anonymous):

Not a problem!

OpenStudy (anonymous):

so if m=4 what do i do from there?

OpenStudy (anonymous):

The formula for the tangent line of a function \(f(x)\) at the point \((a, f(a))\) is given by:\[ y-f(a) = f'(a)(x-a) \]

OpenStudy (anonymous):

The formula for the tangent line of a function \(f(x)\) at the point \((a, f(a))\) is given by:\[ y-f(a) = f'(a)(x-a) \] Given that \(a=4\) and \(f(a)=(16-3)(16+1)\)\[ y-(16-3)(16+1) = f'(4)(x-4) \]